RHIC Run-11 Setup Bumps on the Road

RHIC Run-11 Setup Bumps on the Road
 

    Jan. 8

    §         WFG failed to ramp bo6-tv9 to its requested current.

    Jan. 9

    §         Micardo and SVD orbit correction did not work.

    §         BLM was unmasked during snake on setup, but probably unknown to MCR. As the snake current was brought on gradually, no quench happened.

    Jan. 10

    §         OptiCalc slow, it took 99% CPU time.

    Jan. 13

    §         Yellow injection kicker timing hardware issue. It took us three shifts to sort it out.

    Jan. 14

    §         BLM first turn trigger problem and log problem.

    §         L20 tripped many times.

    Jan. 15

    §         LINAC Mod8 problem from time to time.

    Jan. 16

    §         The orbit feedback target orbits does not include the snake bumps.

    §         LINAC mode8 problem continues

    Jan. 18

    §         111 bunch injection caused blue2 polarimeter Si 4 to have high leakage current.

    Jan. 19

    §         Main quad glitch on the ramp caused significant beam loss, sometimes beam abort..

    §         Vacuum valves near polarimeters closed when switching target in yellow.

    §         A bad BPM (yo8-bh4) wrong polarity) was put back online which caused confusion.

    Jan. 21

    §         Orbit feedback defined its separation bumps but did not use design one. Fixed by Mei and Al.

    §         Upstream polarimeter (outside ring) do not work.

    Jan. 22

    §         When correctors did not start ramp when ramp started, operation failed to abort beam but rely on BLM pulling the permit. It does not work all the time.

    Jan. 25

    §         LINAC Mod8 is not available for seven hours.

    Jan. 27

    §         No rotator ramp handy. The one produced in-situ had wrong IP4 DX magnet setting which caused DX magnet quench.

    §         Tape sequencer is not working properly when dealing with two ramps. Copy stone does not work as expected.

    §         Collimators were at wrong settings, which caused beam dumps.

    Jan. 28

    §         Orbit feedback pushed vertical orbit off by 1.6mm which caused heavy losses at collimators and beam induced quench. Six hours recover time.

    §         Due to the bug in the RhicLossThreshold, it can save/change setting without warning. The BLM events were all masked and did not pull the permit for the ramp which caused beam induced quench.

    Jan. 29

    §         Vacuum vavles at polarimeters were closed during polarization measurements and dumped beam.

    Jan. 31

    §         corrector P/S problem in RHIC holds the operation off for eight hours.

    Feb. 5

    §         Start seen emittance growth in blue.

    Feb. 7

    §         AGS RF cavity water leak, need ring access.

    §         AGS warm snake power supply failures, five hours lost.

    §         RHIC blue dump kicker noise problem.

    Feb. 8

    §         RHIC blue dump kicker noise problem.

    Feb. 9

    §         RHIC blue dump kicker noise problem.

    Feb. 11

    §         RHIC blue dump kicker noise problem(was fixed at the end of day).

    Feb. 14

    §         RHIC cryo power distribution system failure. Lost 14 hours.

    Feb. 16

    §         RHIC PASS system communication problem. Lost 5 hours.

    §         RHIC LLRF problem. Lost 7 hours.

    Feb. 17

    §         RHIC LLRF problem affected rotator setup efficiency and beam intensity in RHIC (limited to less than 1.2*10^11).